Please help answering all parts
ignment Problem 3.36 A proton and an electron are separated by 8.5 x 10-10 m (a) Calculate the magnitude of the electric force that the proton exerts on the electron on electron (b) Calculate the magnitude of the electric force that the electron exerts on the proton. on proton/ = 30
Answer #1

Problem 3.36:

Given,

The distance by which the electron and proton are separated is = 8.5 * 10-10 m

(a)

The magnitude of charge on electron and proton , e = 1.602 *10-19 C

The magnitude of the electric force, |F| = k * e2 / d2

= (9 * 109 * 1.602 * 10-19)2 / ( 8.5 * 10-10)2

= 0.3196 * 10-9

   = 3.196 * 10-10 N

(b)

The magnitude of charge on electron and proton , e = 1.602 *10-19 C

The magnitude of the electric force, |F| = k * e2 / d2

= (9 * 109 * 1.602 * 10-19)2 / ( 8.5 * 10-10)2

= 0.3196 * 10-9

   = 3.196 * 10-10 N